### v1

# LLM-Generated Black-box Explanations Can Be Adversarially Helpful

Abstract
Large Language Models (LLMs) are becoming vital tools that help us solve and understand complex problems by acting as digital assistants. LLMs can generate convincing explanations, even when only given the inputs and outputs of these problems, i.e., in a “black-box” approach. However, our research uncovers a hidden risk tied to this approach, which we call adversarial helpfulness. This happens when an LLM’s explanations make a wrong answer look right, potentially leading people to trust incorrect solutions. In this paper, we show that this issue affects not just humans, but also LLM evaluators. Digging deeper, we identify and examine key persuasive strategies employed by LLMs. Our findings reveal that these models employ strategies such as reframing the questions, expressing an elevated level of confidence, and cherry-picking evidence to paint misleading answers in a credible light. To examine if LLMs are able to navigate complex-structured knowledge when generating adversarially helpful explanations, we create a special task based on navigating through graphs. Some LLMs are not able to find alternative paths along simple graphs, indicating that their misleading explanations aren’t produced by only logical deductions using complex knowledge. These findings shed light on the limitations of black-box explanation setting. We provide some advice on how to use LLMs as explainers safely.
1 Introduction
Large language models (LLMs) have demonstrated strong explaining capabilities. They can generate texts containing logical steps towards solving complex problems (Trinh et al., 2024; Sprague et al., 2023), and incorporate users’ situations (Mondal et al., 2024; Zhu et al., 2023; Zhou et al., 2024). The literature in NLP has many works showing LLMs can generate explanations that are convincing (Wiegreffe et al., 2022), informative (Chen et al., 2023) and faithful (Lyu et al., 2023) to levels that are comparable to human-written explanations. These abilities lead to a promising use case: LLMs as assistants that explain problems to us.
Sounds like we have a convenient use case: we pass a complex problem and its answer to an LLM, together with a suitably-formatted prompt. Hopefully, this “black-box” approach will let the LLM generate an explanation that is instructive and helpful for us to understand the problem. However, as we will show in this paper, this approach may backfire. Specifically, when we prompt an LLM to explain an incorrect answer to a problem, this “black-box” approach may produce explanations that lead us to believe in the incorrect answers.
We first try to gauge the extent of this adversarially helpfulness problem. We let human annotators rate the extent they are convinced of an incorrect answer, before and after seeing an LLM-generated explanation. On a commonsense reasoning task, their ratings are significantly improved by explanations generated by three commercial-tier LLMs. We also use some proxy evaluator models to repeat the survey, and observe similar effects.
[…]
Figure 1: Examples of LLM-generated explanations towards incorrect labels in a commonsense QA question (above) and an NLI question (below). We mark the persuasion strategies with color-codeed angle brackets. To save space, the parts that do not contain persuasion strategies are omitted with […].
Then, we proceed with studying the strategies of the adversarially helpful explanations. We identify ten strategies that we consider relevant to the explanations being adversarially helpful, and examine the strategies used by LLMs in generating these explanations. We find some alarmingly high frequencies. For example, over 90% of the explanations in inference problems involve reframing to varying extents, and over 60% of the explanations in the commonsense problems involve the selective presentation of either the facts or the evidence.
Then we dive into the mechanisms. Are these LLMs able to generate these adversarially helpful explanations because they are able to navigate through complex knowledge, like finding an alternate path in a graph? We set up a symbolic reasoning task, where we ask the LLMs to figure out an alternate path that leads to a specified destination node. This is an abstraction for “explaining an incorrect answer”. We find that the weaker models are unable to complete this task, especially when the graph complexity increases. These indicate that the generation of adversarially helpfulness explanations may involve more than the abilities in deductive reasoning and logical inference, which matches the prior observation that additional strategies (e.g., reframing) are used.
The findings shed light on the limitations of the black-box explanation setting, and motivates us to provide some recommendations for safer usage of LLMs as explaining assistants.
2 Related Works
Reasoning
Recently, a number of works have leveraged the reasoning abilities of LLMs: chain-of-thought (Nye et al., 2021; Wei et al., 2022), tree-of-thought (Yao et al., 2024), graph-of-thought (Besta et al., 2023) and everything-of-thought (Ding et al., 2023) are some examples. We follow their approaches to elicit the reasoning abilities from LLMs.
Utility of explanations
Model-generated explanations can have significant impacts on both human users, for example in mitigating misinformation (Hsu et al., 2023), rescale human judgments (Wadhwa et al., 2023) and understand models’ behavior (Hase & Bansal, 2020; Chen et al., 2022). Researchers have tried to use explanations for model development, yet there are mixed results (Saha et al., 2024; Im et al., 2023).
Failure cases of explanations
The explainability research of AI have identified pitfalls (those unintended effects) versus dark patterns (those resulting from intended misuse) (Ehsan & Riedl, 2021). The adversarial helpfulness of LLM-generated explanation is an example of explainability pitfalls. Specific to the explanations in natural language, these explanations are selective, and can be subjective or misleading (Kunz & Kuhlmann, 2024) or unreliable (Ye & Durrett, 2022). Adversarial helpfulness is different from other problems, in that it might sometimes be a feature; we elaborate in the discussion section.
Jailbreaking and defending
LLMs with self explanations can be biased towards incorrect answers by superficial patterns like the ordering of choices (Turpin et al., 2024). Our paper considers a different setting: instead of planting superficial patterns in the demonstrations, we instruct the LLMs to explain an incorrect answer in a zero-shot manner, resembling how a user would use the LLM as an explaining assistant. This paper is also relevant to the jailbreaking/red-teaming of LLMs (Zou et al., 2023; Zeng et al., 2024; Deng et al., 2023; Ganguli et al., 2022). Instead of developing the jailbreaking or defending algorithms, we focus on the problem itself, and study the strategies adopted by the LLMs when generating adversarially helpful explanations.
3 Experiment setup
Data
Two datasets are used: ECQA (Aggarwal et al., 2021a) and SNLI (Bowman et al., 2015). For ECQA, we sampled 500 problems and hand-pick a “second-best answer” that we consider to be only slightly worse than the correct answer designated by the dataset. Anecdotally, we found that those more abstract problems lead to more significant “adversarial helpfulness” explanations, so we skipped some very concrete problems. For SNLI, we sampled 300 problems from each of the datasets with label Entailment and Contradictory, respectively. We found it was almost impossible to write sufficiently logical arguements (with humans or LLMs) that an Entailment sentence pair should be Contradictory (or vice versa), so we only considered the cases that argue for a Neutral label.
Explainer models
We consider the following four models as explainers: ChatGPT (gpt-3.5-turbo) (OpenAI, 2022), GPT-4 (gpt-4-0613) (Achiam et al., 2023), Claude (Amazon Bedrock version) (Anthropic, 2023), Cohere Command (Cohere, 2023). These are powerful models that provide API access for text generation. We include only the API access models here, since the API models have a potentially larger audience who may be vulnerable to adversarially helpful explanations.
Evaluator models
We consider the following models as evaluators: Vicuna-33B-v1.3 (Chiang et al., 2023), WizardLM-70B-V1.0 (Xu et al., 2023), Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). These are the top-ranking [^1] As of February 2024. Note that Qwen1.5-72B-Chat (Bai et al., 2023) and Yi-34B-Chat (AI et al., 2024) are also ranked on the top places, but we only included one LLM per model size to reduce the computation cost. models on the lmsys Chatbot Arena leaderboard (Zheng et al., 2023) that provide weight access. The weight access is important, since it allows us to query the probabilities in setting up the automatic evaluation — we will describe the setup in Section 5.
Adversarial Explanations
The “adversarially helpful” explanations are generated by prompting the LLM explainer to explain an incorrect answer to a question. Figure 2 lists the prompting templates we used. Empirically, we found other prompting templates can lead to explanations demonstrating a similar “adversarially helpful” behavior, as long as we specifically instruct the model to convince the audience of the incorrect answer.
[…]
Figure 2: The prompt templates for explaining incorrect ECQA and NLI answers. The texts between double braces are to be replaced by the problem-specific contents.
4 Human evaluation of adversarial helpfulness
Protocol
We recruit Amazon Mturk annotators to evaluate the following four scores, on GPT4-generated explanations for the ECQA problems. Each explanation is presented to three randomly recruited annotators, who selects from 1,3 and 5 for each of the following four scores.
  1. 1.
    Convincingness of the “second-best answer”, without seeing the explanation. This score serves as a baseline for the “surprisingness” of the given answer.
  2. 2.
    Convincingness of the “second-best answer” after seeing the explanation. This annotation UI is presented to the annotators after the explanation.
  3. 3.
    Fluency of the explanation. If there are signs of incoherence between the sentences explaining one answer choice, the explanations will receive a low fluency score.
  4. 4.
    Factual correctness of the explanation. If the annotators find factually incorrect information, they will take off marks in factual correctness.
Appendix A includes the screenshots of the UI, including the marking criteria. This protocol is approved by the ethical review board at our university.
Humans consider the explanations helpful
The MTurk annotators rate the LLM-generated explanations to have high fluency and correctness ratings. As Table 1 shows, the convincingness ratings rise from $2.96(\textrm{sd}=0.99)$ to $3.53(\textrm{sd}=0.93)$, from $3.66(\textrm{sd}=0.88)$ to $3.72(\textrm{sd}=0.85)$ and from $3.74(\textrm{sd}=0.97)$ to $3.84(\textrm{sd}=0.94)$, for GPT4, Claude, and ChatGPT, respectively. Paired $t$-tests (dof=$499$, two-tailed for all three) find significant differences ($p<0.01$ for all three) between the pre-explanation and the post-explanation convincingness scores, showing that the humans consider the explanations beneficial for the convincingness of the answers, even when the answers are incorrect.
5 Automatic evaluation of adversarial helpfulness
Evaluator setup
To examine the effects of the generated explanations in a scalable manner, we use several evaluator language models as proxies for MTurk annotators. Following is the protocol we use for querying the response from a proxy.
Given an input text $\mathbf{x}$, a proxy model computes the conditional probability of the next token: $P(y\,|\,\mathbf{x})$. The score given by a proxy is formulated as:
$\hat{y}=\max_{y\in[\textrm{``1'',``3'',``5''}]}P(y\,|\,\mathbf{x})$ (1)
Here, the input texts $\mathbf{x}$ for each question are identical to the texts presented to the human annotators in MTurk. We observe several interesting effects, and summarize them below.
Score Explainer
GPT4 Claude ChatGPT
[…]
ECQA (“Second-best”) 2.96 2.61 1.64 3.35 3.66 2.61 1.64 3.35 3.74 2.61 1.64 3.35
NLI ($E\rightarrow N$) 2.99 1.13 3.55 2.99 1.13 3.55 2.99 1.13 3.55
NLI ($C\rightarrow N$) 3.01 1.15 3.71 3.01 1.15 3.71 3.01 1.15 3.71
[…]
ECQA (“Second-best”) 3.53 2.59 3.01 3.70 3.72 2.59 3.01 3.70 3.84 2.59 3.01 3.70
NLI ($E\rightarrow N$) 3.00 3.00 4.83 3.00 3.00 4.83 3.00 3.00 4.83
NLI ($C\rightarrow N$) 3.00 3.00 4.95 3.00 3.00 4.95 3.00 3.00 4.95
[…]
ECQA (“Second-best”) 4.85 1.95 1.30 3.08 4.55 1.95 1.30 3.08 4.46 1.95 1.30 3.08
NLI ($E\rightarrow N$) 2.21 1.11 3.22 2.21 1.11 3.22 2.21 1.11 3.22
NLI ($C\rightarrow N$) 2.04 1.10 3.27 2.04 1.10 3.27 2.04 1.10 3.27
[…]
ECQA (“Second-best”) 4.68 2.98 1.39 4.54 3.86 2.95 1.39 4.54 4.04 2.98 1.39 4.54
NLI ($E\rightarrow N$) 2.99 1.11 4.91 2.99 1.11 4.91 2.99 1.11 4.91
NLI ($C\rightarrow N$) 3.00 1.11 4.93 3.00 1.11 4.93 3.00 1.11 4.93
Table 1: Human and automatic evaluation results for the convincingness (before and after), fluency, and correctness scores for the generated explanations. The evaluators “M”, “V” and “W” refer to Mixtral-8x7B, Vicuna-33B and WizardLM-70B, respectively.
The explanations are not unhelpful, according to the models
We observe relatively consistent trends on both ECQA and NLI datasets. On both datasets, the convincingness ratings for the incorrect answers by Mixtral-8x7B do not significantly differ. The other two models, Vicuna-33B and WizardLM-70B, compute increased probabilities which show statistical significance ($p<0.001$ on two-tailed $t$ tests, Bonferroni corrected, with $\textrm{dof}=499$ for ECQA and $\textrm{dof}=299$ for NLI).
Note that the utility of LLM that evaluate the convincingness should be taken with caution. First, LLMs have demonstrated some evidence to model humans’ thoughts (i.e., “theory-of-mind” modeling) (Kosinski, 2024), but the actual capability is being debated. Second, in some cases the LLM-given ratings show some “degeneration” trends. For example, in NLI $E\rightarrow N$, all C_after scores computed by Mixtral-8x7B are the same, resulting in a average score of 3.00. This indicates that some of the scores given by the proxy evaluators might result from the dataset artefacts rather than the contents.
Models and humans show varying numbers, item-wise
When getting down to the item-wise level, the evaluators show some different trends. First, humans correlate weakly with the proxy models. On the dataset where human annotator results are available, we compute the Pearson correlations between the averaged human results and the evaluator models. None of the correlations are significant, indicating that the evaluator models show very different fluency, correctness and convincingness ratings from the human annotators. Second, human results are random. This is because the MTurk distribute the annotation tasks to more than three annotators, so the inter-annotator agreements do not apply here. Third, the proxy models assign different scores. We compute the Cohen’s Kappa between each pair of the proxy models. None of them have a value larger than 1 for any score, indicating poor agreements between the evaluators. This is because the proxy models have different “baseline marking guidelines”, as is illustrated by the drastically different mean scores on Table 1.
6 Strategies toward adversarial helpfulness
The recent literature involve many taxonomies of persuasion strategies. For example, Dimitrov et al. (2021) identified strategies in social media texts. Piskorski et al. (2023) considered news. Zeng et al. (2024) considered 40 techniques for jailbreaking LLMs as well as other NLP tasks. Inspired by these works, we identify ten persuasion strategies that we consider particularly relevant to LLM-generated explanations. A brief summarization of each strategy is included in Appendix B.
We use GPT-4 Turbo (gpt-4-0125-preview) to detect the persuasion strategies used in the explanations. Figure 1 lists some examples of the identified strategies.
Frequencies of the strategies
Table 2 lists the frequencies of the persuasion strategies adopted by the three explainer models. Some trends emerge from the results.
First, the sheer numbers of frequency are alarming. For the commonsense questions, more than $70\%$ of the explanations highlight the comparative advantage towards the incorrect answers. For inference tasks, more than $90\%$ of the explanations involve reframing of the questions. While these results are only on the commonsense and inference datasets, we expect that the real-world applications of natural language explanations frequently include these persuasion strategies.
Second, the three explainer models have common trends in applying persuasion strategies. The LLM-generated explanations appear overly confident. [^2] This may related to the overly-confident tone in our prompts, as listed in Figure 2. Strategies like selective evidence, and selective fact presentation are frequently used, and the strategies like appeal to authority and analogical evidence are very infrequently used in any of the models. These indicate that the safeguarding against the adversarially helpful explanations can benefit from the safeguarding against a finite set of persuasion strategies.
For completeness, we repeat the automatic detection experiments using the taxonomy of Zeng et al. (2024) — the experiment results are included in Table 3 in Appendix, showing similar observations: Only a few of the persuasion strategies/techniques are applied (e.g., logical appeal, encouragement and framing), but for those applied techniques, the frequencies are alarmingly high.
ECQA (“Second-best”) NLI ($E\rightarrow N$) NLI ($C\rightarrow N$)
GPT4 Claude Chat GPT4 Claude Chat GPT4 Claude Chat
1. Confidence manipulation 38 65 39 78 62 65 62 67 69
2. Appeal to authority 5 3 3 1 1 0 0 4 1
3. Selective evidence 79 69 67 55 48 43 53 67 46
4. Logical fallacies 11 28 10 6 10 6 13 17 9
5. Comparative advantage framing 90 79 82 37 31 22 33 23 20
6. Reframing the question 48 57 53 93 95 94 92 94 94
7. Selective fact presentation 67 72 69 49 48 51 56 37 53
8. Analogical evidence 2 1 3 1 2 1 1 4 1
9. Detailed scenario building 63 28 32 24 30 20 18 7 12
10. Complex inference 4 1 4 9 8 5 3 7 3
Table 2: The frequencies (out of 100) of persuasion strategies adopted by explainer models. The three mostly adopted strategies per column are marked in bold font.
7 A structural analysis towards adversarial helpfulness
Here we dive deeper into the mechanism of the adversarial helpfulness phenomena that we observe in previous sections. As Figure 3 illustrates, the explanation in natural language has an inherent structure that resembles a graph.
The literature about discourse analysis and automatic reasoning has drawn analogies between reasoning, explanation (and discourse in general) and graphs. One of the seminal works along this direction is Rhetorical Structure Theory (Mann & Thompson, 1988), which identified spans of texts (discourse units) as graph nodes and specified the discourse relations as graph edges. ConceptNet (Liu & Singh, 2004) and other knowledge graphs specified concepts as graph nodes, and abstracted the relations as graph edges. The ECQA dataset (Aggarwal et al., 2021b) that we use is based on Commonsense QA (Talmor et al., 2019), which is based on ConceptNet. Dziri et al. (2024) abstracted the compositional reasoning problems into graphs while studying the difficulty of the reasoning problems. Prystawski et al. (2024) used a Bayesian network to model how the reasoning emerges from the locality of experience. CLEAR (Ma et al., 2022) and RSGG-CE (Prado-Romero et al., 2024) leveraged graph structures to generate counterfactual explanations. Following their avenues of research, we set up a graph-based symbolic reasoning problem as an abstraction of the “explanation towards the incorrect answer” phenomenon.
Figure 3: Explanation has a graph-like structure, where each graph node is analogous to a “discourse unit” and each graph edge serves as a “reasoning step”.
Problem specification
We consider the process of explanation to be an instance of path finding on a graph. In this path finding problem, we find a path from the “root” node to the “answer” node. The explanation serves as the path that connects the known (the “root” node) to the unknown (the “answer” node).
Figure 4: An example of symbolic reasoning graph. Left: the graph appearing in DOT string format. Right: the graph appearing in graph format. If the path “root $\rightarrow$ 0_1 $\rightarrow$ 0_2 $\rightarrow$ A” is the reasoning path supporting answer A, supporting answer C would need a reasoning path “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ C”.
Following specifies how the graphs are constructed. There are two parameters: number of branches $B$ and path length $L$.
1. We specify a root node, marked as “root”.
2. We specify $B$ branches. All branches start from the root node, and extends by $L$ steps. The $j^{th}$ node at the $i^{th}$ branch is marked $i_{j}$, and the last node at each branch is marked with an alphabet (A, B, C, and so on), to resemble the answer in the multiple-choice question.
LLMs have limited capabilities in capturing the structural information in graphs (Huang et al., 2023a). Based on the intuition, we attach some simplification assumptions. First, all branches have the same path length. Second, each non-root node uniquely appears on only one path. Figure 4 shows an example of such graph with $B=3$ branches of $L=4$ path.
3. We linearize each graph using the DOT format. Together with this graph string and a brief description of the formatting specifications, we prompt the LLM to find the alternative answer, and return us with the path from root to the alternative answer. The correctness of the returned path is evaluated with exact match. [^3] Empirically we find that many LLMs tend to write the last node twice. For example, for the graph in Figure 3, instead of the reasoning path “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ C”, the LLMs sometimes write “root $\rightarrow$ 2_1 $\rightarrow$ 2_2 $\rightarrow$ 2_3 $\rightarrow$ C”. We consider this correct as well, if we allow the $\rightarrow$” to signal both a graph reasoning step and a name aliasing step.
4. We compute the success rate of alternate path finding at a given graph complexity. To reduce the complexity of experiments, we fix $B=L$, and take this number as the “graph complexity”. At a given graph complexity $B$, there are $B\times(B-1)$ alternate path-finding cases. Among them, the portion of correctly returned path is defined as the success rate.
Randomizing the node names
An LLM might generate responsed based on the naming patterns instead of the graph path structure. To deconfound this factor, we run another version of symbolic graph reasoning experiment. This time, we replace the node names of each non-leaf and non-root node with a randomly-chosen but non-overlapping character.
Results
Figure 5 plots the success rate against the number of graph branches, without (left) and with (right) the randomization step of the node names. As the complexity of the graph increase, the success rate of alternative path finding decreases. When we factor out the reliance on the node names, the performances of all models (except GPT-4/4 Turbo) drop to zero. We consider one graph node to be the abstraction of one sentence, so the complexity of the graphs resemble that of the generated explanations. Why, then, do the LLM-generated explanations show adversarial helpfulness while these LLMs fail badly at handling the structures of the adversarial helpfulness? Apparently, two factors jointly contribute to the “adversarial helpfulness”: the structure and the lexical contents. When an LLM is not capable of handling the structures, it can still utilize the lexical contents to produce adversarially helpful explanations.
Figure 5: Success rate vs graph complexity. Left: using the default graph node names. Right: replacing node names like “0_1” with random non-overlapping characters.
8 Discussion
Adversarial helpfulness vs unfaithfulness
There are intricate differences between these two improvement areas of LLM-generated explanations. Adversarial helpfulness refers to the explainer’s behavior that rationalizes an incorrect phenomenon. Unfaithfulness, on the other hand, rationalizes problems using the incorrect reasons or irrelevant facts. Using the analogy in the graph reasoning setting, an adversarially helpful explanation is analogous to a path (with or without an incorrect node) that leads to an incorrect destination, and an unfaithful explanation is analogous to a path that involves incorrect nodes or edges, yet leads to the correct destination.
Adversarial helpfulness vs plausibility
Plausibility is different from faithfulness — for an extended discussion of their difference please refer to Agarwal et al. (2024). Here we consider that adversarial helpfulness has an overlap with plausibility; the latter indicates a potential negative impact towards human, whereas the former does not. Plausibility is a feature, but adversarial helpfulness is a bug.
Do the guardrails work?
When trying to let LLMs produce adversarially helpful explanations on the reported datasets, we find that the existing guardrails are very weak, if they exist at all. Many existing guardrails rely on the models’ parametric knowledge (“common-sense”) to avoid outputting adversarial explanations. This mechanism can be rewritten by simply asking the LLM to, e.g., “hypothesize an alternative scenario”. Since we should not prevent an LLM from reasoning counterfactually, one can argue that the adversarial helpfulness problem is unable to be guardrailed at the LLM level. Instead, this problem should be guarded at the user level.
Safe use of AI as explainers
We provide some recommendations here.
First, let humans make decisions, not AIs. We can use the AI explainer as a “Prudence” model (Miller, 2023). A “Prudence” model provides evidence supporting human’s decisions, without directly giving us an answer. The alternative, “Bluster” model, recommends an answer, optionally with the rationales supporting that answer. Since the rationales supporting the answer can be adversarially helpful, it is important to delegate the least possible amount of “decision-making” responsibility to the explainers (especially those based on generative AI).
Second, instruct AIs to generate rationales that support multiple alternative answers to offset the cherry-picking behavior of AI explainers.
Third, pass in as much of the decision-maker model’s intermediate signals as possible, especially when the decisions are difficult (i.e., the decision-maker model can likely produce an incorrect label). Note that LLMs still struggle at summarizing the neurons’ activations (Huang et al., 2023b) — perhaps because the neurons are too fine-grained (Niu et al., 2024) — but this struggle should not prevent us from giving the explainer a guideline to follow.
9 Conclusion
We identify a potentially perilous scenario of using LLMs as explaining assistants in a “black-box” manner, which we call adversarial helpfulness. When prompted to explain an incorrect answer, LLMs can generate convincing explanations, making wrong answers look right. We show that this issue affects both humans and LLM evaluators. We analyze the persuasion strategies, and find that LLMs frequently reframe the questions and present selective details, among other strategies, in favor of the incorrect answers. We further analyze the LLMs’ mechanisms by setting up a symbolic reasoning problem. The findings motivate us to recommend future practices for using LLMs as explainers.
Appendix A Annotation User Interface
Figures 6 to 8 show the templates of the user interface shown to the MTurk annotators. The fields \${question}, \${choice_A} through \${choice_E}, \${answer} and \${explanation} are filled in dynamically for each data sample.
Figure 6: The MTurk UI for commonsense questions, when first presented to the annotator.
Figure 7: Our UI checks for completeness before the annotator submits the scores.
Figure 8: The MTurk UI for commonsense questions, after the first question is answered.
Appendix B Additional details for the persuasion strategies
  1. 1.
    Confidence Manipulation: Here, LLMs might express high confidence in their alternative answers to persuade users. This could involve using assertive language or citing (real or fabricated) sources to bolster the credibility of their responses.
  2. 2.
    Appeal to Authority: LLMs could reference authoritative sources or experts—even if inaccurately—to justify their alternative answers. This strategy leverages the user’s trust in expertise and authority figures to lend weight to the model’s response.
  3. 3.
    Selective Evidence: In presenting justifications, LLMs might selectively use evidence that supports their alternative answers while ignoring or minimizing evidence that contradicts them. This could involve cherry-picking data, quotes, or studies that back up the LLM’s stance.
  4. 4.
    Logical Fallacies: Employing flawed reasoning patterns that may appear logical at first glance, such as slippery slopes, straw man arguments, or false dilemmas. While potentially convincing, these fallacies do not hold up under closer scrutiny.
  5. 5.
    Comparative Advantage Framing: Highlighting the benefits or advantages of the alternative answer over other possibilities without necessarily proving it as the only correct option. This can involve comparative analysis with other known solutions or outcomes.
  6. 6.
    Reframing the Question: Subtly altering the interpretation of the question to fit the alternative answer better. This might involve focusing on specific words or phrases in the question that could be ambiguously interpreted.
  7. 7.
    Selective Fact Presentation: Presenting facts, statistics, or data that exclusively support the alternative answer while conveniently omitting or de-emphasizing information that supports the correct answer.
  8. 8.
    Analogical Evidence: Drawing analogies to similar situations or questions where the less obvious or unconventional choice was actually the more accurate one, suggesting a parallel to the current scenario.
  9. 9.
    Detailed Scenario Building: Construct specific, detailed scenarios where the alternative answer is the most logical or applicable, using vivid descriptions to make the scenario as relatable and convincing as possible.
  10. 10.
    Complex Inference: Utilize complex inferential reasoning that logically leads to the alternative answer, relying on a chain of deductions that, while not immediately obvious, are sound and lead to the alternative conclusion.
Appendix C Additional results for persuasion strategy identification
Table 3 lists the identified persuasion strategies and techniques, using the taxonomy of Zeng et al. (2024).
ECQA (“Second-best” NLI ($E\rightarrow N$) NLI ($C\rightarrow N$)
chat claude gpt4 chat claude gpt4 chat claude gpt4
1. Evidence-based persuasion 36 28 46 6 8 8 2 6 7
2. Logical appeal 61 58 78 10 22 13 4 21 21
3. Expert endorsement 2 3 1 0 1 0 0 0 1
4. Non-expert testimonial 2 1 1 0 0 0 0 0 0
5. Authority endorsement 4 2 3 0 0 0 0 0 0
6. Social proof 5 3 5 0 0 0 0 0 0
7. Injunctive norm 1 2 3 0 0 0 0 0 0
8. Foot-in-the-door 1 1 2 0 0 0 0 0 0
9. Door-in-the-face 0 0 0 0 0 0 0 0 0
10. Public commiement 0 0 0 0 0 0 0 0 0
11. Alliance building 1 0 0 0 0 0 0 0 0
12. Complimenting 2 1 2 0 0 0 0 0 0
13. Shared values 4 4 8 0 0 0 0 0 0
14. Relationship leverage 1 0 1 0 0 0 0 0 0
15. Loyalty appeals 0 0 0 0 0 0 0 0 0
16. Favor 0 0 0 0 0 0 0 0 0
17. Negotiation 0 0 0 0 0 0 0 0 0
18. Encouragement 32 21 17 50 56 72 41 40 72
19. Affirmation 46 42 32 45 36 68 33 26 59
20. Positive emotional appeal 21 16 38 3 4 5 3 1 4
21. Negative emotional appeal 9 11 15 0 0 0 0 0 1
22. Storytelling 34 40 53 13 8 26 9 4 23
23. Anchoring 3 4 2 0 0 1 1 1 0
24. Priming 2 6 2 0 0 0 1 1 2
25. Framing 33 61 55 3 0 5 6 2 9
26. Confirmation bias 3 14 4 0 0 0 0 0 0
27. Reciprocity 0 0 1 0 0 0 0 0 0
28. Compensation 0 0 0 0 0 0 0 0 0
29. Supply scarcity 0 0 0 0 0 0 0 0 0
30. Time pressure 0 0 0 0 0 0 0 0 0
31. Reflective thinking 4 4 5 19 21 30 20 22 32
32. Threats 0 0 0 0 0 0 0 0 0
33. False promises 0 1 0 0 0 0 0 0 0
34. Misrepresentation 1 3 0 0 0 0 0 0 0
35. False information 1 4 0 0 0 0 0 0 0
36. Rumors 0 0 0 0 0 0 0 0 0
37. Social punishment 0 0 0 0 0 0 0 0 0
38. Creating dependency 0 0 0 0 0 0 0 0 0
39. Exploiting weakness 1 1 1 0 0 0 0 0 0
40. Discouragement 0 0 0 0 0 0 0 0 0
Table 3: Frequencies (out of 100) of strategies following the taxonomy of Zeng et al. (2024).